Journal of the Association for Computing Machinery, Vol. 42, No. 4, July 1995, pp. 908–918.

A Constant-Time Optimal Parallel String-Matching Algorithm

ZVI GALIL

Abstract. Given a pattern string, we describe a way to preprocess it. We design a CRCW-PRAM constant-time optimal parallel algorithm for finding all occurrences of the (preprocessed) pattern in any given text.

Categories and Subject Descriptors. F.2.2 [computations on discrete structures]

General Terms. Algorithms, Theory

Additional Key words and Phrases. Constant time, CRCW-PRAM, lion hunting, optimal parallel algorithm, period, string matching

Copyright 1995 by ACM, Inc. Permission to make digital or hard copies of part or all of this work for personal or classroom use is granted without fee provided that copies are not made or distributed for profit or direct commercial advantage and that copies bear this notice and the full citation on the first page. Copyrights for components of this work owned by others than ACM must be honored. Abstracting with credit is permitted. To copy otherwise, to republish, to post on servers, or to redistribute to lists, requires prior specific permission and/or a fee. Request permissions from Publications Dept, ACM Inc., fax +1 (212) 869-0481, or permissions@acm.org.